# Eigenspectrum Analysis of Neural Networks without Aspect Ratio Bias

### v1

@@ C.5 Detailed PL_Alpha_Hill Distribution
Figure 10: Comparing the distribution of PL Alpha Hill of DPOT-Tiny and DPOT-Small in different fine-tuning methods and data ratios.
@@
Figure 11: Comparing the distribution of PL Alpha Hill of LLaMA-7B and LLaMA-13B in different layer-wise strategies and sparsity ratios (shown as ”SR”). The pruning method is SparseGPT.
@@ C.8 LLM Pruning Stability Analysis
The experiment results from  (Yin et al., 2023) show that layer-wise pruning LLMs at high sparsity is a challenging optimization problem. If sparsity ratios are not properly allocated, the performance of the model can become very unstable. Therefore, here we aim to demonstrate that the performance improvements observed in our experiments are due to more accurate layer-wise analysis, rather than accidental factors such as random seeds.
@@ E.1 Relationship with Matrix Shape
The importance of recognizing that the numerical value of the same property measured in different network layers can naturally have different scales, and thus should not be directly compared without proper normalization. In some real-world application scenarios, such as in LLM Pruning  (Lu et al., 2024; Liu et al., 2025a), Model Merging  (Yang et al., 2024; Zhu et al., 2025), and LLM Reinforcement Learning (Wang et al., 2025, 2024; Chen et al., 2024), this factor need to be considered. Indeed, similar considerations regarding the different shapes of weight matrices in each layer have previously led to impactful theoretical and practical advancements. For example, works such as Tensor Programs IV  (Yang and Hu, 2021) and Tensor Programs V  (Yang et al., 2022) have explored how layer dimensions affect the optimal scaling of initializations and learning rates. Tensor Programs IV introduced the Maximal Update Parametrization (µP) to ensure feature learning by carefully choosing parameter scaling rules based on these dimensions, addressing how standard parametrizations might otherwise collapse to kernel regimes in the infinite-width limit. Building on this, Tensor Programs V demonstrated that µP can enable zero-shot hyperparameter transfer, where hyperparameters tuned on smaller models remain optimal for significantly larger ones. In this context, ”shape” awareness pertains to designing parametrizations that maintain stable training dynamics.
Our work also concerned with the influence of matrix shape, addresses a different challenge: the bias introduced by the aspect ratio (number of rows versus columns) of an individual weight matrix on the measurement of its ESD. We demonstrate that varying aspect ratios can artificially stretch or compress the ESD. This distortion confounds the interpretation of heavy-tailed (HT) metrics. To mitigate this measurement bias, we propose FARMS to technique analyze average ESD of submatrices sampled at a fixed aspect ratio. This approach provides a normalized HT metric, enabling more reliable comparisons of such spectral diagnostics across different layers within a given network. Thus, while and focus on shape-aware parametrization for training stability and transfer, our contribution lies in a shape-aware analysis technique (FARMS) aimed at correcting measurement bias in spectral diagnostics.
@@ E.3 Does this approach introduce additional bias?
Further justification for this perspective comes from recent work  (Wang et al., 2023; Kothapalli et al., 2024) that aims to theoretically quantify heavy-tailedness. These studies interpret heavy-tailedness as the accumulation and evolution of feature spikes in the ESD that align with the teacher model’s features. Critically, these feature spikes are characterized as being approximately rank-one updates to the original matrix Because a rank-one component inherently covers the whole matrix, sampling a submatrix will, with high probability, capture that rank-one component. Therefore, this subsampling process is unlikely to miss the feature spikes, which are identified by previous work as the cause of the heavy-tail structure. We believe this provides substantial evidence that FARMS can preserve important spectral information, specifically as measured by these feature spikes in the ESD.
@@ Appendix F Hyperparameter Adjustment
CAL 0.05, 0.1, 0.15 - 78.23$\pm$0.087
TB(no LS) 0.1 (0.6, 1.4) 78.76$\pm$0.111
ResNet 18 LS+TB 0.1 (0.2, 1.8) 79.31$\pm$0.180
Ours(no LS) 0.1 (0.1, 1.9) 79.49$\pm$0.080
LS+Ours 0.1 (0.1, 1.9) 79.53$\pm$0.177
CAL 0.05, 0.1, 0.15 - 78.99$\pm$0.137
TB(no LS) 0.1 (0.5, 1.5) 79.64$\pm$0.029
ResNet 34 LS+TB 0.1 (0.3, 1.7) 80.00$\pm$0.090
Ours(no LS) 0.1 (0.2, 1.8) 80.17$\pm$0.213
LS+Ours 0.1 (0.3, 1.7) 80.20$\pm$0.221
CAL 0.025, 0.05, 0.1 - 74.30$\pm$0.078
TB(no LS) 0.05 (0.6, 1.4) 74.43$\pm$0.158
VGG 16 LS+TB 0.05 (0.3, 1.7) 75.19$\pm$0.131
Ours(no LS) 0.05 (0.2, 1.8) 75.36$\pm$0.118
LS+Ours 0.05 (0.2, 1.8) 75.15$\pm$0.247
CAL 0.025, 0.05, 0.1 - 73.11$\pm$0.113
TB(no LS) 0.05 (0.6, 1.4) 73.22$\pm$0.277
VGG 19 LS+TB 0.05 (0.2, 1.8) 74.19$\pm$0.159
Ours(no LS) 0.05 (0.2, 1.8) 74.28$\pm$0.392
LS+Ours 0.05 (0.4, 1.6) 73.99$\pm$0.300

### v2

@@ C.5 Detailed PL_Alpha_Hill Distribution
Figure 10: Comparing the distribution of PL_Alpha_Hill of DPOT-Tiny and DPOT-Small in different fine-tuning methods and data ratios.
@@
Figure 11: Comparing the distribution of PL_Alpha_Hill of LLaMA-7B and LLaMA-13B in different layer-wise strategies and sparsity ratios (shown as ”SR”). The pruning method is SparseGPT.
@@ C.8 LLM Pruning Stability Analysis
The experiment results from  (Yin et al., 2024) show that layer-wise pruning LLMs at high sparsity is a challenging optimization problem. If sparsity ratios are not properly allocated, the performance of the model can become very unstable. Therefore, here we aim to demonstrate that the performance improvements observed in our experiments are due to more accurate layer-wise analysis, rather than accidental factors such as random seeds.
@@ E.1 Relationship with Matrix Shape
One should recognize that the numerical value of the same property measured in different network layers can naturally have different scales, and thus should not be directly compared without proper normalization. In some real-world application scenarios, such as in LLM Pruning  (Lu et al., 2024; Liu et al., 2025a), this factor need to be considered. When the network size changes, hyperparameters also need to be changed accordingly. For example, works such as Tensor Programs IV  (Yang and Hu, 2021) and Tensor Programs V  (Yang et al., 2021) have explored how layer sizes affect the optimal scaling of weight initializations and learning rates. Tensor Programs IV introduced the Maximal Update Parametrization (µP) to ensure feature learning by carefully choosing parameter scaling rules based on these sizes, addressing how existing parametrizations might otherwise collapse to kernel regimes. Building on this, Tensor Programs V demonstrated that µP can enable zero-shot hyperparameter transfer, where hyperparameters tuned on smaller models remain optimal for significantly larger ones. In this context, “shape” awareness pertains to designing size-dependent hyperparameters that maintain stable and effective (non-kernel) training dynamics.
Nevertheless, our work is fine-grained in that we consider the aspect ratios of different layers (number of rows versus columns). We demonstrate that varying aspect ratios can artificially stretch or compress the ESD. This distortion confounds the interpretation of heavy-tailed (HT) metrics. To mitigate this measurement bias, we propose FARMS to analyze the average ESD of submatrices sampled at a fixed aspect ratio. This approach provides a normalized HT metric, enabling more reliable comparisons of such spectral diagnostics across different layers within a given network. Thus, while focusing on shape-aware parametrization for training stability and transfer, our contribution lies in a shape-aware analysis technique (FARMS) aimed at correcting measurement bias in spectral diagnostics.
@@ E.3 Does this approach introduce additional bias?
Further justification for this perspective comes from recent work  (Wang et al., 2023; Kothapalli et al., 2025) that aims to theoretically quantify heavy-tailedness. These studies interpret heavy-tailedness as the accumulation and evolution of feature spikes in the ESD that align with the teacher model’s features. Critically, these feature spikes are characterized as being approximately rank-one updates to the original matrix Because a rank-one component inherently covers the whole matrix, sampling a submatrix will, with high probability, capture that rank-one component. Therefore, this subsampling process is unlikely to miss the feature spikes, which are identified by previous work as the cause of the heavy-tail structure. We believe this provides substantial evidence that FARMS can preserve important spectral information, specifically as measured by these feature spikes in the ESD.
@@ Appendix F Hyperparameter Adjustment
[…]